# Network Intrusion Datasets: A Survey, Limitations, and Recommendations

### v1

@@ 7 Related Work
This section compares the presented survey with other works discussing network intrusion detection datasets. While data-specific NID surveys (this paper) are rare, general NID surveys often include datasets as well. Naturally, they typically lack detailed discussion on data-specific topics yet still provide a relevant reference point for many readers. We thus compare our work to those focused explicitly on NID data and five general surveys with data-related sections, summarized in Table 4.
@@
Based on related work analysis, we conclude that our paper:
@@
The closest study to ours, and its inspiration in many aspects, is the work by Ring et al. [13]. Considered one of the first surveys on NID data, it formalized data properties, compared existing datasets, and provided data-related recommendations. However, as it was published in 2019, it does not include newer datasets. Our study builds on this work by modifying some established properties (e.g., removing the data balance property, as NID data are naturally imbalanced) and incorporating newly released datasets. We also add sections on data limitations and trends to address recent findings and emphasize data quality, aiming to provide a holistic view of NID data.
@@
Although Hindy et al. [44] cover aspects similar to this paper, they primarily focus on the network threat landscape and its relation to 21 NID datasets (until 2018). The study further discusses NIDS algorithms’ and datasets’ popularity, threat taxonomy, IDS limitations, and recommendations for future research. In contrast, our study includes more datasets, compares them by several properties rather than by attacks, and provides a more detailed discussion of NID data-related aspects.
Yang et al. [178] conduct an SLR on the NID domain as a whole, highlighting popular ML-based methods and datasets for benchmarking. Although not explicitly data-focused, the study provides the largest published list of datasets (52), compares them, and analyzes their popularity, with KDD'99 and NSL-KDD coming as the most popular. In contrast, our SLR focuses specifically on the data, listing more datasets and elaborating on data-specific topics absent in Yang et al.’s work.
@@ 8 Conclusions
Our analysis of contemporary state-of-the-art NIDS research revealed a shift in benchmark dataset popularity: DARPA-based datasets are no longer the most prevalent. While CIC-IDS2017 is now frequently used, researchers still prefer to collect their own data, most of which are not publicly shared. This trend, while understandable due to the limitations of existing datasets, hinders reproducibility and validation. We emphasize the importance of following the best practices presented in this paper, as well as sharing the custom data to improve the quality and reliability of future NIDS research.
@@
Given NID domain-specific properties and resulting data limitations discussed in this paper, we state that creating a perfect dataset is infeasible. We consider data timeliness as a key issue, as research relying on static datasets will always lag behind evolving attacks and network trends. A step towards mitigating the issue could be achieved by adopting continuously updated datasets, as discussed in the next section.
9 Future Domain Directions
9.1 Data Generation
The research branch of data generation focuses on generating timely and realistic data – either by collecting real-world traces or improving traffic emulation techniques.
Collecting timely real-world data: One of the key NID domain limitations is the lack of realistic real-world patterns. As our survey revealed, only one in five datasets contains real traffic, with this ratio decreasing in recent years. This limitation impacts NIDS generalization, leading to overfitting to unrealistic traffic patterns and reduced performance in operational environments. For this reason, more data with contemporary real-world network patterns are required. While datasets like CTU-13 [55], Appraise H2020 [155], and LITNET [127] incorporate real traffic, they remain static and thus quickly become outdated as attack patterns evolve.
Improving simulated normal traffic: Although obtaining real-world traffic is usually considered ideal, it is often difficult or impossible to capture due to privacy concerns, lack of desired attack events, or operational constraints. For this reason, realistic traffic emulation remains crucial for constructing diverse and representative NID datasets. While attack traffic can be emulated via existing tools used by real-world attackers, simulating realistic benign traffic remains the main challenge.
To enhance traffic realism and robustness, future research should focus on closing the gap between simulated (emulated and synthetic) benign traffic and its real-world counterpart. Although models like Generative Adversarial Networks (GANs) [185] and Bayesian Networks (BNs) [52] have already been explored, investigating and comparing novel generative methods like diffusion models [228] for benign network traffic synthesis offer promising research directions. Furthermore, the synthetic traffic must respect real-world constraints, making it important to formalize domain-specific conditions for synthetic traffic validity (e.g., network flows cannot have zero bytes).
Despite their benefits, reproducibility and extensibility have received little attention from the community. Therefore, most virtualized testbed datasets lack environment setup files, configuration details, or information on customizing the environment. Although AIT-LDSv2.0 [43] emphasized reproducibility, its environment customizability is limited. Future research should thus focus on proposing standardized methods, frameworks, and templates for fully replicable and extensible NID data creation in virtualized and cloud environments to enable dataset validation, modification, customization, and extension.
In addition to previous scenarios, Hindy et al. [44] and Keersmaeker et al. [222] point out that many attacks and IoT environments are not covered within public NID datasets at all. For these reasons, future research should prioritize covering specific scenarios and environments, along with ensuring robust coverage of emerging threats to streamline NIDS research in these areas. However, without continually captured and extensible datasets is made, new datasets will remain susceptible to getting outdated due to the ever-changing nature of NID.
9.2 Data Processing
Exploring data granularities for intrusion detection: As shown in our survey, most datasets are flow-based, while nearly half miss packet data. Flow-based NID has become the de facto standard for NID research. Similarly, many studies explored the selection of optimal flow-based features [230, 231, 232], but little research has been done on their effectiveness across different intrusion detection scenarios.
9.3 Data Validation
Formalizing frameworks for quality evaluation: While some metrics for NID data quality evaluation exist [203, 103], their primary goal is to compare d atasets rather than assess their intrinsic quality (i.e., realism, correctness, and usability). Furthermore, their interpretation is often ambiguous due to the lack of usage instructions. These facts make evaluating and selecting suitable NID datasets difficult, leading to frequent utilization of flawed datasets that undermine research reliability.
In order to improve NID data evaluation, future research should formalize methods and develop frameworks for evaluating NID data validity and quality, i.e., ensure datasets are both realistic and error-free. Current data validation heavily relies on manual analysis, as seen in studies on CIC-IDS2017 and CSE-CIC-IDS2018 issues [109, 45, 164], with Engelen et al. [227] recommending manual analysis by domain experts before public data release. This validation type is time-consuming and dependent on the availability of experts, posing additional challenges for data creators. Therefore, research on NID data quality should also integrate expert knowledge into (semi-)automated validation methods. Standardized approaches would benefit data creators to validate their datasets and assist users in selecting suitable data, enhancing NID research relevance, trustworthiness, and explainability.
9.4 Data Publishing
Formalizing data documentations: While reviewing the metadata of surveyed datasets, we observed that each was different. Although documentation of all scientifically-backed datasets follows a standard paper structure, some offer detailed data descriptions while others mention the data only briefly. This leads to under-documented datasets, reducing their credibility and usability. In addition, some datasets lack scientific backing or any documentation altogether.
@@ A.1 Data Search and Filtering Process
In order to make our research replicable, we specify the way how we identified and filtered the datasets included in the survey (Figure 8). The process was as follows:
Figure 8: The diagram depicts the steps of searching, filtering, and selecting the relevant studies throughout the systematic literature review process. Practically, we obtained the objects of interest from three sources: other survey papers, our own search via paper titles and keywords, and the analysis of backward and forward references. In order to make the reference analysis feasible, we performed it after the initial filtering. Finally, all the discovered sources went through quality assurance to include only relevant datasets for others.
@@
    Listing known datasets: We extracted known datasets from older network intrusion surveys listed in Table 4, resulting in 302 datasets (with duplicates).
@@
    Own literature search: We conducted a search with the following string: network (intrusion $\cup$ anomaly $\cup$ outlier $\cup$ attack $\cup$ threat) (dataset $\cup$ data set). The publication year was limited to 2023, so papers published since 2024 are not included. Matching on title or author keywords, we used the following search engines and databases:
@@
    • Google Scholar: 600 results [^24] Since Google Scholar returns thousands of results, we reviewed only the first 20 search pages (200 results) for all years, along with the first ten pages for 2020, 2021, 2022, and 2023 (400 results) to find new datasets not covered in previous surveys. We are aware that this process is not exactly replicable. However, omitting it would cause over a dozen datasets to stay undiscovered.
@@
    Relevance filtering: After collecting 302 + 1477 items from the previous steps, we excluded duplicates. We selected only dataset-specific papers, as most articles only used existing datasets. Filtering was done via title, abstract, and full-text analysis using the following criteria:
@@
    • Presents an IDS method using an existing data,
@@
    Backward and forward reference search: In order to obtain the most comprehensive contemporary dataset list, we additionally reviewed papers citing (forward search) and cited by (backward search) the papers obtained from the third step using Google Scholar. These actions were limited to the papers from the past five years (2018-2023). After filtering the relevant references using the criteria from the previous step, this resulted in 25 additional papers, bringing the number to 138.
@@
    Quality assurance: Lastly, we aimed to ensure the quality of the selected papers. Although datasets backed by a scientific publication should be preferred, a non-negligible portion of the data is only presented via informal means (e.g., a website), so including them is also desirable.
@@
    • Collected as an auxiliary material to other research not primarily focused on the data,
    • Heavily pre-processed data preventing meaningful analysis and sound conclusions,
    • Data not publicly available.
After applying the quality assurance checks, we were left with 89 datasets for a final review.
@@ A.2 Data Popularity Determination
In order to aid in the dataset selection, we also analyze dataset popularity in contemporary NIDS research. For this purpose, we focus on network intrusion/anomaly detection from the cybersecurity Tier 1 and Tier 2 conferences, according to the MLSec group list [^25] MLSec is a research group focused on the intersection of cybersecurity and machine learning at Technical University (TU) Berlin. They list top-tier cybersecurity conferences at https://mlsec.org/topnotch/. Similar rankings are also maintained elsewhere, e.g., Guofei Gu: https://people.engr.tamu.edu/guofei/sec_conf_stat.htm.. In particular, we surveyed the following 12 conferences: S&P (Oakland), CCS, Security, NDSS, ACSAC, AsiaCCS, CSF, ESORICS, EuroS&P, PETS, RAID, and DIMVA, in addition to the MLSec list.
Aiming to maximize our findings’ relevance, we analyzed only papers from the past four years, i.e., 2020 – 2023. This was performed by manually reviewing conference proceedings and conducting full-text analyses to identify used datasets.

### v2

@@ 8 Related Work
This section compares the presented survey with other works discussing network intrusion detection datasets. While data-specific NID surveys (this paper) are rare, general NID surveys often include datasets as well. Naturally, they typically lack detailed discussion on data-specific topics yet still provide a relevant reference point for many readers. We thus compare our work to those focused explicitly on NID data and five general surveys with data-related sections, summarized in Table 5. Based on the related work analysis, we conclude that our paper:
@@
The closest study to ours, and its inspiration in many aspects, is the work by Ring et al. [14]. Considered one of the first surveys on NID data, it formalized data properties, compared existing datasets, and provided data-related recommendations. However, since it was published in 2019, it does not include newer datasets. Our study builds on this work by refining some established properties (e.g., removing the data balance property, as NID data are naturally imbalanced) and incorporating newly released datasets. We also introduce sections on data limitations and trends to address recent findings and emphasize data quality, aiming to provide a holistic view of NID data.
@@
Although Hindy et al. [46] cover aspects similar to this paper, they primarily focus on the network threat landscape and its relation to 21 NID datasets (until 2018). The study further discusses NIDS algorithms’ and datasets’ popularity, threat taxonomy, IDS limitations, and recommendations for future research. In contrast, our study includes more datasets, compares them by their properties rather than by attacks, and provides a more detailed discussion of NID data-related aspects.
Yang et al. [180] conducted an SLR on the NID domain as a whole, highlighting popular ML-based methods and datasets for benchmarking. Although not explicitly data-focused, the study presents the largest published dataset list at the time (52), compares them, and analyzes their popularity, with KDD'99 and NSL-KDD coming as the most popular. In contrast, our SLR specifically focuses on the data, listing more datasets and elaborating on data-specific topics absent in Yang et al.’s work.
@@ 9 Conclusions
Our analysis of contemporary state-of-the-art NIDS research revealed a shift in benchmark dataset popularity: DARPA-based datasets are no longer the most prevalent. While CIC-IDS2017 is now frequently used, researchers still prefer to collect their own data, most of which are not publicly shared. This trend, while understandable due to the limitations of existing datasets, hinders reproducibility and validation. We emphasize the importance of following the best practices discussed in this paper, as well as sharing the data to improve the quality and reliability of future NIDS research.
@@
Given the NID domain-specific properties and resulting data limitations outlined in this paper, we state that creating a perfect dataset is infeasible. We consider data timeliness as a key issue, as research relying on static datasets will inevitably lag behind ever-evolving attacks and network trends. A step towards mitigating the issue could be achieved by adopting continuously updated datasets, as discussed in Section 7.
@@ A.1 Data Search and Filtering Process
In order to ensure the replicability of our research, we outline the process used to identify and filter the datasets included in the survey (Figure 8). The steps were as follows:
Figure 8: The diagram illustrates the steps of searching, filtering, and selecting relevant studies during the systematic literature review process. Specifically, we identified the objects of interest from three sources: existing survey papers, our own search via paper titles and keywords, and an analysis of backward and forward references. To ensure the feasibility of the reference analysis, it was conducted after the initial filtering. Finally, all identified sources underwent a quality assurance process to include only relevant datasets.
@@
    Listing known datasets: We extracted known datasets from existing network intrusion surveys listed in Table 5, resulting in 302 datasets (with duplicates).
@@
    Own literature search: We conducted a search with the following query: network (intrusion $\cup$ anomaly $\cup$ outlier $\cup$ attack $\cup$ threat) (dataset $\cup$ data set). The publication year was limited to 2023, so papers published since 2024 are not included. Matching on titles or author keywords, we used the following search engines and databases:
@@
    • Google Scholar: 600 results [^24] Since Google Scholar returns thousands of results, we reviewed only the first 20 search pages (200 results) for all years, along with the first ten pages for 2020, 2021, 2022, and 2023 (400 results) to identify new datasets not covered in previous surveys. We acknowledge that this process is not entirely replicable. However, omitting it would leave over a dozen datasets undiscovered.
@@
    Relevance filtering: After collecting 302 + 1477 items from the previous steps, we excluded duplicates. Next, we selected only dataset-specific resources, as most articles from the previous step only used existing datasets. This filtering was performed based on title, abstract, and full-text analysis using the following criteria:
@@
    • Presents an IDS method using existing data,
@@
    Backward and forward reference search: Aiming to compile the most comprehensive contemporary dataset list, we also reviewed papers that cited (forward search) and were cited by (backward search) those obtained in the third step using Google Scholar. This process was limited to papers from the past five years (2018-2023). Applying the same filtering criteria as in the previous step yielded 25 additional papers, bringing the total to 138.
@@
    Quality assurance: Lastly, we aimed to ensure the quality of the selected papers. While datasets backed by scientific publications are preferable, a non-negligible portion of the data is only available via informal sources (e.g., websites), so including them is also desirable.
@@
    • Was collected as auxiliary material to other research not primarily focused on data,
    • Is heavily pre-processed, thus preventing meaningful analysis and sound conclusions,
    • Is not publicly available.
Failing to reject all criteria results in the exclusion of a dataset. After their application, 89 datasets remained for a final review.
@@ A.2 Data Popularity Determination
In order to aid in the dataset selection, we also analyze dataset popularity in contemporary NIDS research. For this purpose, we focus on network intrusion/anomaly detection from Tier 1 and Tier 2 cybersecurity conferences, according to the MLSec group list [^25] MLSec is a research group at Technical University (TU) Berlin focused on the intersection of cybersecurity and machine learning. They list top-tier cybersecurity conferences at https://mlsec.org/topnotch/. Similar rankings are also maintained by others, e.g., Guofei Gu: https://people.engr.tamu.edu/guofei/sec_conf_stat.htm.. In particular, we surveyed the following 12 conferences: S&P (Oakland), CCS, Security, NDSS, ACSAC, AsiaCCS, CSF, ESORICS, EuroS&P, PETS, RAID, and DIMVA, in addition to the MLSec list.
Aiming to maximize the relevance of our findings, we analyzed only papers from the past four years (2020-2023). This was done by manually reviewing conference proceedings and conducting full-text analyses to identify used datasets.